OpenStudy (anonymous):

Write the slope -intercept equation for the line that passes through (1,-9) and (3,-1)? would my answer Y=4x-13

jimthompson5910 (jim_thompson5910):

You are 100% correct, nice job

OpenStudy (anonymous):

thanks!

jimthompson5910 (jim_thompson5910):

You can verify your answer by plugging in each point Checking (1,-9) y=4x-13 -9=4(1)-13 -9=4-13 -9=-9 ... true -------------------- Checking (3,-1) y=4x-13 -1=4(3)-13 -1=12-13 -1=-1 ... true So both points lie on the line which confirms the answer.